# The KSBA moduli space of stable log Calabi–Yau surfaces

### 23 Feb 2024 v1

@@ 6.4.1. The polarized mirror family to an open Kulikov degeneration
Let $\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. As reviewed in Proposition 3.36, the contraction $\mathcal{X}\rightarrow\overline{\mathcal{X}}$ is a relative Mori dream space. The corresponding Mori fan $MF(\mathcal{X}/\overline{\mathcal{X}})$ is a complete fan in $\mathrm{Pic}(\mathcal{X}/\overline{\mathcal{X}})_{\mathbb{R}}$, whose maximal dimensional cones are indexed by the finitely many $\mathbb{Q}$-factorial birational models $f^{\prime}:\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$ of $\mathcal{X}$ over $\overline{\mathcal{X}}$, and are given by $\mathrm{Nef}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})+\langle\mathrm{Ex}(f^{\prime})\rangle$, where $\mathrm{Nef}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})$ is the relative nef cone of $\mathcal{X}^{\prime}$ over $\overline{\mathcal{X}}$, and $\langle\mathrm{Ex}(f^{\prime})\rangle$ is the non-negative span of the $f^{\prime}$-exceptional divisors.
@@
The proof is analogous to the proof of Theorem [49, Theorem 2.12]. ∎
@@
Conversely, if $\mathrm{Nef}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})$ and $\mathrm{Nef}(\mathcal{X}^{\prime\prime}/\overline{\mathcal{X}})$ are equivalent, then, by Lemma 6.18, $\mathrm{Nef}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})$ and $\mathrm{Nef}(\mathcal{X}^{\prime\prime}/\overline{\mathcal{X}})$ are both contained in a maximal cone $C$ of $\mathrm{MovSec}(\mathcal{X}/\overline{\mathcal{X}})$. As $C$ is convex, there exists a path contained in $C$, starting in $\mathrm{Nef}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})$, ending $\mathrm{Nef}(\mathcal{X}^{\prime\prime}/\overline{\mathcal{X}})$, and only intersecting cones of $MF(\mathcal{X}/\overline{\mathcal{X}})$ of codimension zero or one. Therefore, it is enough to prove the result when $\mathrm{Nef}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})$ and $\mathrm{Nef}(\mathcal{X}^{\prime\prime}/\overline{\mathcal{X}})$ share a common codimension one face, that is, when $\mathcal{X}^{\prime}$ and $\mathcal{X}^{\prime\prime}$ are related by an elementary flop, which by Lemma 6.1 is either a M1 or a M2 flop. By Lemma 6.5, $\mathcal{X}^{\prime}$ and $\mathcal{X}^{\prime\prime}$ would not be equivalent if they were related by an M2 flop, and so one concludes that $\mathcal{X}^{\prime}$ and $\mathcal{X}^{\prime\prime}$ are in fact related by an M1 flop. ∎
@@
Let $\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}}$ be the polarized family as in (6.5) with the ample line bundle $\mathcal{L}_{\mathcal{X}}^{\mathrm{msec}}$. We define the theta divisor in the linear system of $\mathcal{L}_{\mathcal{X}}^{\mathrm{msec}}$ as
$\mathcal{C}_{\mathcal{X}}^{\mathrm{msec}}:=\left\{\sum_{p\in P_{\mathbb{Z}}}\vartheta_{p}=0\right\}\in|\mathcal{L}_{\mathcal{X}}^{\mathrm{msec}}|\,.$ (6.6)
We denote by $\mathcal{C}_{X}$ the restriction of $\mathcal{C}_{\mathcal{X}}^{\mathrm{msec}}$ to $\mathcal{Y}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$.
@@
6.4.2. The divisor $\mathcal{D}_{\mathcal{X}}$
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration, with tropicalization $(C(P),C(\mathscr{P}))$, and let $r\in\partial C(P)_{\mathbb{Z}}$. Then, for every $p,q\in C(P)_{\mathbb{Z}}$ and $\beta\in NE(\mathcal{X})$, we have
@@
    $N_{pqr}^{\beta}(\mathcal{X})=1$ if there exists $i\in\mathbb{Z}/N\mathbb{Z}$ such that $p,q,r\in C([v_{i},v_{i+1}])$, $r=p+q$ inside $C([v_{i},v_{i+1}])$, and $\beta=\varphi_{i}(p)+\varphi_{i}(q)-\varphi_{i}(r)$.
@@
    otherwise, $N_{pqr}^{\beta}(\mathcal{X})=0$.
@@
If $p=0$ or $q=0$, the result follows from [47, Theorem 4.3]. Similarly, if $r=0$, the result follows immediately from Definition 5.3. So, from now on, we assume that $p,q,r$ are all non-zero. Denote by $Z_{p}$, $Z_{q}$, and $Z_{r}$ the strata of $\mathcal{X}_{0}$ corresponding to the smallest cones of $C(P)$ containing $p$, $q$, and $r$ respectively. By Definition 5.3, a punctured map contributing to $N_{pqr}^{\beta}(\mathcal{X})$ has two marked points mapping to $Z_{p}$, $Z_{q}$, and one marked point mapping to a general point of $Z_{r}$. The assumption $r\in(\partial C(P)_{\mathbb{Z}})$ implies that $Z_{r}$ is either a 0-surface or a 1-surface. But it follows from Definitions 3.2-3.4 that there are no compact curve passing through a general point of a 0-surface, and that the only compact curve passing through a general point of a 1-surface is a $\mathbb{P}^{1}$-fiber of the corresponding generically $\mathbb{P}^{1}$-fibration. Hence, we have $N_{pqr}^{\beta}(\mathcal{X})=0$, unless $p,q,r$ are all contained in a common cone $C([v_{i},v_{i+1}])$.
In the latter case, the contributing punctured maps have image contained in a chain of $\mathbb{P}^{1}$-fibers of intersecting $1$-surfaces, and so the invariants $N_{pqr}^{\beta}(\mathcal{X})$ are equal to the corresponding invariants for a toric Calabi–Yau 3-fold whose fan contains $C([v_{i},v_{i+1}])$ as a boundary face. By [47], the invariants $N_{pqr}^{\beta}(\mathcal{X})$ can be computed in terms of broken lines in the canonical scattering diagram. In the toric case, the broken lines are never bending and only gets contributions from the kinks of the PL function $\varphi_{i}$ while crossing the rays $\mathbb{R}_{\geq 0}v_{i,j}$ – see [16, §2], and so the statement of Lemma 6.22 i) follows. ∎
@@
Let $\pi\colon\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration, with tropicalization $(C(P),C(\mathscr{P}))$. Then, $\mathcal{D}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$ has toric irreducible components and is a flat family of divisors in the fibers of $\mathcal{Y}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$. Moreover, the general fiber of $\mathcal{D}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$ is a cycle of smooth rational curves $D_{i,i+1}$ in one-to-one correspondence with the intervals $[v_{i},v_{i+1}]\subset\partial P$, and the central fiber $D_{0_{\mathcal{X}}}$ over $0_{\mathcal{X}}\in S_{X}$ is a cycle of smooth rational curves $D_{v}$ in one to one correspondence with the boundary integral points $v\in P_{\mathbb{Z}}$. In addition, the theta functions $\vartheta_{p}$ with $p\in\partial C(P)_{\mathbb{Z}}$ naturally restrict to toric monomials on the irreducible components of $\mathcal{D}_{\mathcal{X}}$.
@@
for $i\in\mathbb{Z}/N\mathbb{Z}$, and that $\mathcal{D}_{\mathcal{X},i,i+1}\rightarrow S_{\mathcal{X}}$ is the toric Mumford degeneration defined by the polyhedral decomposition of $C([v_{i},v_{i+1}])$ given in Equation (6.7), and the PL function $\varphi_{i}$. In particular, the general fiber of $\mathcal{D}_{\mathcal{X},i,i+1}\rightarrow S_{\mathcal{X}}$ is $\mathbb{P}^{1}$, and its central fiber over $0_{\mathcal{X}}$ is a chain of $\mathbb{P}^{1}$-components indexed by the integral points $v\in[v_{i},v_{i+1}]\cap P_{\mathbb{Z}}$. This implies Lemma 6.24. ∎
@@ 6.5. Torus action and extension over the bogus cones
Let $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}},\mathcal{L}_{\mathcal{X}}^{\mathrm{msec}})\rightarrow S_{\mathcal{X}}^{\mathrm{msec}}$ be the polarized family as in Equation (6.5), $\mathcal{C}_{\mathcal{X}}^{\mathrm{msec}}$ the theta divisor as in Definition 6.21, and $\mathcal{D}_{\mathcal{X}}^{\mathrm{sec}}$ the divisor as in Lemma 6.25. In this section, we describe how to canonically extend the family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}},\mathcal{D}^{\mathrm{msec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}_{\mathcal{X}}^{\mathrm{msec}})$ defined over $S_{\mathcal{X}}^{\mathrm{sec}}$, to a family over the toric variety with fan given by the “secondary fan” defined as follows.
@@
We denote from now on by $S^{\mathrm{sec}}_{\mathcal{X}}$ the proper toric variety whose fan is the secondary fan $\mathrm{Sec}(\mathcal{X}/\overline{\mathcal{X}})$. In what follows, we show that the family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}},\mathcal{D}^{\mathrm{msec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}_{\mathcal{X}}^{\mathrm{msec}})$ extends to a family $(\mathcal{Y}^{\mathrm{sec}}_{\mathcal{X}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}}+\epsilon\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}})$ over $S^{\mathrm{sec}}_{\mathcal{X}}$. To do this, we first describe a natural torus action on $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}},\mathcal{D}^{\mathrm{msec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}_{\mathcal{X}}^{\mathrm{msec}})$.
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. Denote by $T_{0}$ the torus $\mathrm{Spec}\,\mathbb{C}[\mathbb{Z}^{P_{\mathbb{Z}}}]$. Note that $\mathbb{Z}^{P_{\mathbb{Z}}}$ can be naturally identified with the group of divisors of $\mathcal{X}$ supported on $\mathcal{X}_{0}$, since the points of $P_{\mathbb{Z}}$ are in a natural one-to-one correspondence with the irreducible components of $\mathcal{X}_{0}$. In particular, for every divisor $F$ supported on $\mathcal{X}_{0}$, we have a corresponding one-parameter subgroup $\mathbb{C}^{\star}\subset T_{0}$. There is a natural action of $T_{0}$ on $S_{\mathcal{X}}=\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}/\overline{\mathcal{X}})]$ such that, for every divisor $F$ supported on $\mathcal{X}_{0}$, the corresponding one-parameter subgroup acts with weights $F\cdot\beta$ on the monomials $t^{\beta}\in S_{\mathcal{X}}$.
@@
This follows immediately from Lemma 6.7– see also [47, Construction 1.24] in the general context of the intrinsic mirror construction. ∎
The action of the torus $T_{0}$ on the affine toric varieties $S_{\mathcal{X}^{\prime}}$ corresponding to the projective crepant resolutions $\mathcal{X}^{\prime}\rightarrow\overline{\mathcal{X}}$ naturally induces an action of $T_{0}$ on the toric varieties $S^{\mathrm{\mathrm{msec}}}_{\mathcal{X}}$ and $S^{\mathrm{sec}}_{\mathcal{X}}$. It follows from the Equation (6.4) that the actions of $T_{0}$ on the families $\mathcal{Y}_{\mathcal{X}^{\prime}}\to S_{\mathcal{X}^{\prime}}$ given by Lemma 6.27 glue into an action of $T_{0}$ on $\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}}\rightarrow S_{\mathcal{X}}^{\mathrm{msec}}$.
@@
There exists a unique $T_{0}$ equivariant extension of the family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}},\mathcal{D}^{\mathrm{msec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}_{\mathcal{X}}^{\mathrm{msec}})\rightarrow S^{\mathrm{\mathrm{msec}}}_{\mathcal{X}}$ to a family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{sec}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}})\to S^{\mathrm{sec}}_{\mathcal{X}}$.
@@
This follows by the existence of the action of $T_{0}$ on $\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}}\to S_{\mathcal{X}}^{\mathrm{\mathrm{msec}}}$ and [49, Corollary 5.7]. ∎
@@
By Theorem 6.28, the line bundle $\mathcal{L}_{\mathcal{X}}^{\mathrm{msec}}$ on $\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}}$ extends to a line bundle $\mathcal{L}_{\mathcal{X}}^{\mathrm{sec}}$ on $\mathcal{Y}_{\mathcal{X}}^{\mathrm{sec}}$. Moreover, by Lemma 6.27, the theta functions $\vartheta_{p}$ with $p\in P_{\mathbb{Z}}$ extend to sections of $\mathcal{L}_{\mathcal{X}}^{\mathrm{sec}}$. By Definition 6.26, every cone of the secondary fan $\mathrm{Mov}(\mathcal{X}/\overline{\mathcal{X}})$ is of the form $\rho+\sum_{v\in I}[X^{v}]$, where $\rho$ is a cone of $\mathrm{MovSec}(\mathcal{X}/\overline{\mathcal{X}})$ and $I$ is a set of integral points $v\in P_{\mathbb{Z}}$ such that the corresponding irreducible components $X^{v}$ of $\mathcal{X}_{0}$ are either 1-surfaces or 2-surfaces. By [49, §6.4], in restriction to the toric stratum of $S^{\mathrm{sec}}_{\mathcal{X}}$ corresponding to the cone $\rho+\sum_{v\in I}\mathbb{R}_{\geq 0}[X^{v}]$ of $\mathrm{Sec}(\mathcal{X}/\overline{\mathcal{X}})$, the divisor $\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}}$ given by Theorem 6.28 is given by the equation $\sum_{p\in P_{\mathbb{Z}}\setminus I}\vartheta_{p}=0$.

### 30 Mar 2024 v2

@@ 6.4.1. The polarized mirror family to an open Kulikov degeneration
For a quasi-projective open Kulikov degeneration $\mathcal{X}\rightarrow\Delta$, the contraction $\mathcal{X}\rightarrow\overline{\mathcal{X}}$ is a relative Mori dream space by Proposition 3.36. The corresponding Mori fan $MF(\mathcal{X}/\overline{\mathcal{X}})$ is a complete fan in $\mathrm{Pic}(\mathcal{X}/\overline{\mathcal{X}})_{\mathbb{R}}$, whose maximal dimensional cones are indexed by the finitely many $\mathbb{Q}$-factorial birational models $f^{\prime}:\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$ of $\mathcal{X}$ over $\overline{\mathcal{X}}$, and are given by $\mathrm{Nef}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})+\langle\mathrm{Ex}(f^{\prime})\rangle$, where $\mathrm{Nef}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})$ is the relative nef cone of $\mathcal{X}^{\prime}$ over $\overline{\mathcal{X}}$, and $\langle\mathrm{Ex}(f^{\prime})\rangle$ is the non-negative span of the $f^{\prime}$-exceptional divisors.
@@
The proof is analogous to the proof of Theorem [50, Theorem 2.12]. ∎
@@
Conversely, if $\mathrm{Nef}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})$ and $\mathrm{Nef}(\mathcal{X}^{\prime\prime}/\overline{\mathcal{X}})$ are equivalent, then, by Lemma 6.18, $\mathrm{Nef}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})$ and $\mathrm{Nef}(\mathcal{X}^{\prime\prime}/\overline{\mathcal{X}})$ are both contained in a maximal cone $C$ of $\mathrm{MovSec}(\mathcal{X}/\overline{\mathcal{X}})$. As $C$ is convex, there exists a path contained in $C$, starting in $\mathrm{Nef}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})$, ending $\mathrm{Nef}(\mathcal{X}^{\prime\prime}/\overline{\mathcal{X}})$, and only intersecting cones of $MF(\mathcal{X}/\overline{\mathcal{X}})$ of codimension zero or one. Therefore, it is enough to prove the result when $\mathrm{Nef}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})$ and $\mathrm{Nef}(\mathcal{X}^{\prime\prime}/\overline{\mathcal{X}})$ share a common codimension one face, that is, when $\mathcal{X}^{\prime}$ and $\mathcal{X}^{\prime\prime}$ are related by an elementary flop, which by Lemma 6.1 is either a M1 or a M2 flop. By Lemma 6.5, $\mathcal{X}^{\prime}$ and $\mathcal{X}^{\prime\prime}$ would not be equivalent if they were related by an M2 flop, and so we conclude that $\mathcal{X}^{\prime}$ and $\mathcal{X}^{\prime\prime}$ are in fact related by an M1 flop. ∎
@@
Let $\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}}$ be the family as in Equation (6.5) with the ample line bundle $\mathcal{L}_{\mathcal{X}}^{\mathrm{msec}}$. We define the theta divisor in the linear system of $\mathcal{L}_{\mathcal{X}}^{\mathrm{msec}}$ as
$\mathcal{C}_{\mathcal{X}}^{\mathrm{msec}}:=\left\{\sum_{p\in P_{\mathbb{Z}}}\vartheta_{p}=0\right\}\in|\mathcal{L}_{\mathcal{X}}^{\mathrm{msec}}|\,,$ (6.6)
and we denote by $\mathcal{C}_{X}$ the restriction of $\mathcal{C}_{\mathcal{X}}^{\mathrm{msec}}$ to $\mathcal{Y}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$.
@@
6.4.2. The divisor $\mathcal{D}^{\mathrm{msec}}_{\mathcal{X}}$
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration, with tropicalization $(C(P),C(\mathscr{P}))$, and let $r\in\partial C(P)_{\mathbb{Z}}$. Then, for every $p,q\in C(P)_{\mathbb{Z}}$ and $\beta\in NE(\mathcal{X})$, we have either
@@
    $N_{pqr}^{\beta}(\mathcal{X})=1$ if there exists $i\in\mathbb{Z}/N\mathbb{Z}$ such that $p,q,r\in C([v_{i},v_{i+1}])$, $r=p+q$ inside $C([v_{i},v_{i+1}])$, and $\beta=\varphi_{i}(p)+\varphi_{i}(q)-\varphi_{i}(r)$, or
@@
    $N_{pqr}^{\beta}(\mathcal{X})=0$ elsewise.
@@
If $p=0$ or $q=0$, the result follows from [48, Theorem 4.3]. Similarly, if $r=0$, the result follows immediately from Definition 5.3. So, from now on, we assume that $p,q,r$ are all non-zero. Denote by $Z_{p}$, $Z_{q}$, and $Z_{r}$ the strata of $\mathcal{X}_{0}$ corresponding to the smallest cones of $C(P)$ containing $p$, $q$, and $r$ respectively. By Definition 5.3, a punctured map contributing to $N_{pqr}^{\beta}(\mathcal{X})$ has two marked points mapping to $Z_{p}$, $Z_{q}$, and one marked point mapping to a general point of $Z_{r}$. The assumption $r\in(\partial C(P)_{\mathbb{Z}})$ implies that $Z_{r}$ is either a 0-surface or a 1-surface. But it follows from Definitions 3.2-3.4 that there are no compact curve passing through a general point of a 0-surface, and that the only compact curve passing through a general point of a 1-surface is a $\mathbb{P}^{1}$-fiber of the corresponding generically $\mathbb{P}^{1}$-fibration. Hence, we have $N_{pqr}^{\beta}(\mathcal{X})=0$, unless $p,q,r$ are all contained in a common cone $C([v_{i},v_{i+1}])$.
If $p,q,r$ are all contained in a common cone $C([v_{i},v_{i+1}])$, the contributing punctured maps have image contained in a chain of $\mathbb{P}^{1}$-fibers of intersecting $1$-surfaces, and so the invariants $N_{pqr}^{\beta}(\mathcal{X})$ are equal to the corresponding invariants for a toric Calabi–Yau 3-fold whose fan contains $C([v_{i},v_{i+1}])$ as a boundary face. By [48], the invariants $N_{pqr}^{\beta}(\mathcal{X})$ can be computed in terms of broken lines in the canonical scattering diagram. In the toric case, the broken lines are never bending and only gets contributions from the kinks of the PL function $\varphi_{i}$ while crossing the rays $\mathbb{R}_{\geq 0}v_{i,j}$ – see [16, §2], and so the statement of Lemma 6.22 i) follows. ∎
@@
Let $\pi\colon\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration, with tropicalization $(C(P),C(\mathscr{P}))$. Then, $\mathcal{D}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$ has toric irreducible components and is a flat family of divisors in the fibers of $\mathcal{Y}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$. Moreover, the general fiber of $\mathcal{D}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$ is a cycle of smooth rational curves $D_{i,i+1}$ in one-to-one correspondence with the intervals $[v_{i},v_{i+1}]\subset\partial P$, and the central fiber $D_{0_{\mathcal{X}}}$ over $0_{\mathcal{X}}\in S_{X}$ is a cycle of smooth rational curves $D_{v}$ in one-to-one correspondence with the boundary integral points $v\in P_{\mathbb{Z}}$. In addition, the theta functions $\vartheta_{p}$ with $p\in\partial C(P)_{\mathbb{Z}}$ naturally restrict to toric monomials on the irreducible components of $\mathcal{D}_{\mathcal{X}}$.
@@
for $i\in\mathbb{Z}/N\mathbb{Z}$, and that $\mathcal{D}_{\mathcal{X},i,i+1}\rightarrow S_{\mathcal{X}}$ is the Mumford toric degeneration defined by the polyhedral decomposition of $C([v_{i},v_{i+1}])$ given in Equation (6.7), and the PL function $\varphi_{i}$. In particular, the general fiber of $\mathcal{D}_{\mathcal{X},i,i+1}\rightarrow S_{\mathcal{X}}$ is $\mathbb{P}^{1}$, and its central fiber over $0_{\mathcal{X}}$ is a chain of $\mathbb{P}^{1}$-components indexed by the integral points $v\in[v_{i},v_{i+1}]\cap P_{\mathbb{Z}}$. Hence the result follows. ∎
@@ 6.5. Torus action and extension over the bogus cones
Let $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}},\mathcal{L}_{\mathcal{X}}^{\mathrm{msec}})\rightarrow S_{\mathcal{X}}^{\mathrm{msec}}$ be the polarized family as in Equation (6.5), $\mathcal{C}_{\mathcal{X}}^{\mathrm{msec}}$ the theta divisor as in Definition 6.21, and $\mathcal{D}_{\mathcal{X}}^{\mathrm{sec}}$ the divisor as in Lemma 6.25. In this section, we describe how to canonically extend the family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}},\mathcal{D}^{\mathrm{msec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}_{\mathcal{X}}^{\mathrm{msec}})$ defined over $S_{\mathcal{X}}^{\mathrm{msec}}$, to a family over the toric stack with stacky fan given by the stacky “secondary fan” defined as follows.
@@
For every bogus cone $\sigma=\mathrm{Nef}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})+\langle\mathrm{Ex}(f^{\prime})\rangle$, let $M_{\sigma}$ be the sublattice of $\mathrm{Span}(\sigma)\cap\mathrm{Pic}(\mathcal{X}/\overline{\mathcal{X}})$ spanned by $\mathrm{Nef}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})\cap\mathrm{Pic}(\mathcal{X}/\overline{\mathcal{X}})$ and by the classes of irreducible components of $\mathcal{X}_{0}$ contracted by $\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$. Equivalent bogus cones define the same sublattices, and so we have a well-defined sublattice $M_{\sigma}$ for every cone $\sigma$ of the secondary fan. The data of the sublattices $M_{\sigma}$ is a toric stacky data as in [94, Definition 4.1], and so defines a stacky fan $\mathcal{S}ec(\mathcal{X}/\overline{\mathcal{X}})$ with underlying fan $\mathrm{Sec}(\mathcal{X}/\overline{\mathcal{X}})$.
We denote from now on by $S^{\mathrm{sec}}_{\mathcal{X}}$ the proper toric variety whose fan is the secondary fan $\mathrm{Sec}(\mathcal{X}/\overline{\mathcal{X}})$, and by $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$ the proper toric Deligne–Mumford stack with stacky fan $\mathcal{S}ec(\mathcal{X}/\overline{\mathcal{X}})$ – see [94, §4.1] and [38] for the general theory of stacky fans and toric stacks. In what follows, we show that the family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}},\mathcal{D}^{\mathrm{msec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}_{\mathcal{X}}^{\mathrm{msec}})$ extends to a family $(\mathcal{Y}^{\mathrm{sec}}_{\mathcal{X}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}})$ over $\mathcal{S}^{sec}_{\mathcal{X}}$. To do this, we first describe a natural torus action on $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}},\mathcal{D}^{\mathrm{msec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}_{\mathcal{X}}^{\mathrm{msec}})$. Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. Denote by $T_{0}$ the torus $\mathrm{Spec}\,\mathbb{C}[\mathbb{Z}^{P_{\mathbb{Z}}}]$. Note that $\mathbb{Z}^{P_{\mathbb{Z}}}$ can be naturally identified with the group of divisors of $\mathcal{X}$ supported on $\mathcal{X}_{0}$, since the points of $P_{\mathbb{Z}}$ are in a natural one-to-one correspondence with the irreducible components of $\mathcal{X}_{0}$. In particular, for every divisor $F$ supported on $\mathcal{X}_{0}$, we have a corresponding one-parameter subgroup $\mathbb{C}^{\star}\subset T_{0}$. There is a natural action of $T_{0}$ on $S_{\mathcal{X}}=\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}/\overline{\mathcal{X}})]$ such that, for every divisor $F$ supported on $\mathcal{X}_{0}$, the corresponding one-parameter subgroup acts with weights $F\cdot\beta$ on the monomials $t^{\beta}$.
@@
This follows immediately from Lemma 6.7– see also [48, Construction 1.24] in the general context of the intrinsic mirror construction. ∎
The action of the torus $T_{0}$ on the affine toric varieties $S_{\mathcal{X}^{\prime}}$ corresponding to the projective crepant resolutions $\mathcal{X}^{\prime}\rightarrow\overline{\mathcal{X}}$ naturally induces an action of $T_{0}$ on the toric varieties $S^{\mathrm{\mathrm{msec}}}_{\mathcal{X}}$ and $S^{\mathrm{sec}}_{\mathcal{X}}$, and on the toric stack $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$. It follows from the Equation (6.4) that the actions of $T_{0}$ on the families $\mathcal{Y}_{\mathcal{X}^{\prime}}\to S_{\mathcal{X}^{\prime}}$ given by Lemma 6.27 glue into an action of $T_{0}$ on $\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}}\rightarrow S_{\mathcal{X}}^{\mathrm{msec}}$.
@@
There exists a unique $T_{0}$ equivariant extension of the family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}},\mathcal{D}^{\mathrm{msec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}_{\mathcal{X}}^{\mathrm{msec}})\rightarrow S^{\mathrm{\mathrm{msec}}}_{\mathcal{X}}$ to a family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{sec}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}})\to\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$.
@@
This follows by the existence of the action of $T_{0}$ on $\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}}\to S_{\mathcal{X}}^{\mathrm{\mathrm{msec}}}$ and [50, Corollary 5.7]. ∎
@@
By Theorem 6.28, the line bundle $\mathcal{L}_{\mathcal{X}}^{\mathrm{msec}}$ on $\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}}$ extends to a line bundle $\mathcal{L}_{\mathcal{X}}^{\mathrm{sec}}$ on $\mathcal{Y}_{\mathcal{X}}^{\mathrm{sec}}$. Moreover, by Lemma 6.27, the theta functions $\vartheta_{p}$ with $p\in P_{\mathbb{Z}}$ extend to sections of $\mathcal{L}_{\mathcal{X}}^{\mathrm{sec}}$. By Definition 6.26, every cone of the secondary fan $\mathrm{Mov}(\mathcal{X}/\overline{\mathcal{X}})$ is of the form $\rho+\sum_{v\in I}[X^{v}]$, where $\rho$ is a cone of $\mathrm{MovSec}(\mathcal{X}/\overline{\mathcal{X}})$ and $I$ is a set of integral points $v\in P_{\mathbb{Z}}$ such that the corresponding irreducible components $X^{v}$ of $\mathcal{X}_{0}$ are either 1-surfaces or 2-surfaces. By [50, §6.4], in restriction to the toric stratum of $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$ corresponding to the cone $\rho+\sum_{v\in I}\mathbb{R}_{\geq 0}[X^{v}]$ of $\mathrm{Sec}(\mathcal{X}/\overline{\mathcal{X}})$, the divisor $\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}}$ given by Theorem 6.28 is given by the equation $\sum_{p\in P_{\mathbb{Z}}\setminus I}\vartheta_{p}=0$.